A Windlass Drum Has A diameter of 0.8 meter. Show us a Formula and Calculate The length of Rope taken in when the drum turns through 16 Revolutions. Please answer this its very complicated because he did not have a given length of a cylinder windlass drum. please answer this Sir/mam THANK YOU.
Answer by rothauserc(4718)

Each revolution is the circumference of the windlass drum. We know the diameter is .8 meter. Let n be the number of revolutions and note that circumference of a circle is pi * D, then we can write the following formula,
length of rope = n * pi * D where * means multiply, then
length of rope = 16 * pi * .8 = 40.21 meters